OpenStudy (anonymous):

Solve using square root property (4x-1)^2=16

OpenStudy (anonymous):

\[(4x-1)^2=16\] \[4x-1=4\] or \[4x-1=-4\] solve each separately

OpenStudy (anonymous):

[-3/4,5/4]

OpenStudy (anonymous):

yes that would do it

OpenStudy (anonymous):

thank you